PN 30 (explicit) document

PN 30 (implicit from previous page) document No TPIC picture

4 PN 5 (implicit from previous page) Note that IA has nothing to do with PN!

document PN 30 (explicit)

7 PN 30 (implicit from previous page)

8 PN 30 (implicit from previous page) and PN 5 in the middle

9